OpenStudy (anonymous):

Change –5x + 2y = 12 into slope-intercept form.

OpenStudy (anonymous):

y = mx+b -5x + 2y = 12 2y = 5x + 12 y = 5/2x + 6

OpenStudy (anonymous):

Do You Use y=mx+b

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Slope intercept form: \[y=mx+b\